# A Mixed-Order Phase Transition in Continuous-Variable Quantum Networks

### v1

@@ Abstract
Quantum networks (QNs) have been predominantly driven by discrete-variable (DV) architectures. Yet, many optical platforms naturally generate Gaussian states—the common states of continuous-variable (CV) systems, making CV-based QNs an attractive route toward scalable, chip-integrated quantum computation and communication. To bridge the conceptual gap between well-studied DV entanglement percolation theories and their CV counterpart, we introduce a Gaussian-to-Gaussian entanglement distribution scheme that deterministically transports two-mode squeezed vacuum states across large CV networks. Analysis of the scheme’s collective behavior using statistical-physics methods reveals a new form of entanglement percolation—negativity percolation theory (NegPT)—characterized by a bounded entanglement measure called the ratio negativity. We discover that NegPT exhibits a mixed-order phase transition, marked simultaneously by both an abrupt change in global entanglement and a long-range correlation between nodes. This distinctive behavior places CV-based QNs in a new universality class, fundamentally distinct from DV systems. Additionally, the abruptness of this transition introduces a critical vulnerability of CV-based QNs: conventional feedback mechanism becomes inherently unstable near the threshold, highlighting practical implications for stabilizing large-scale CV-based QNs. Our results not only unify statistical models for CV-based entanglement distribution but also uncover previously unexplored critical phenomena unique to CV systems, providing valuable insights and guidelines essential for developing robust, feedback-stabilized QNs.
Table 1: Percolation theories (PT) of entanglement distribution on QNs: DV $\to$ continuous vs. CV $\to$ discontinuous.
Entanglement distribution scheme Percolation theory (measure) Phase transition type
@@
DV Deterministic entanglement transmission (DET) ConPT (Concurrence $c$) Continuous (second order)
CV Gaussian-to-Gaussian (G-G) DET NegPT (Ratio negativity $\chi$) Discontinuous (mixed order)
@@
Quantum information technologies [1] fundamentally rely on our ability to distribute entanglement [2, *JPHH1997], the process of establishing quantum correlations between distant nodes across network architectures, from chips to the internet. This capability empowers applications ranging from quantum computation [4, *Hu2024] to secure quantum communication [6, *HWJP1998, *NCHN2011, *Munro2012]. The collective behavior of entanglement distribution across large-scale systems is often described by the quantum network (QN) model [10]. Within this framework, a QN is represented as a graph where nodes represent spatially separated parties, and links denote bipartite entangled states shared between adjacent nodes. The distribution process itself is executed via protocols based on local operations and classical communication (LOCC), such as entanglement swapping [11, *SVP1999] and entanglement concentration [13, *FM2001], along with more sophisticated schemes [15, *DES2017, *RS2017, *MHD2019], applied across the network’s links to establish long-range correlations.
@@
Investigating how LOCC schemes distribute entanglement across a QN has uncovered fundamental correspondences to percolation theories. The ability to establish long-range entanglement in QN often mirrors the emergence of large-scale connectivity described by percolation theories—the statistical physics framework governing phase transitions on graphs [19, *cross-probab-cft_c92, *netw-percolation_ceah00, *BOCCALETTI2006, *ARENAS2008, *hypergr-percolation_bd23, *hypergr-percolation_lglk23]. This conceptual bridge between entanglement distribution and percolation was first built in 2007, linking a probabilistic entanglement distribution scheme [10] to classical bond percolation [26]. Later, the development of a more effective deterministic entanglement transmission (DET) scheme [27] (where LOCC steps succeed deterministically, eliminating classical randomness unlike earlier probabilistic schemes) led to its mapping onto an essentially distinct variant, concurrence percolation theory (ConPT) [28, *OXSGBJ2022]. This interplay continues with more recent mappings identified between quantum memories and continuum percolation [30], and between entanglement routing and path percolation [31, *path-percolation_mhrk24], underscoring how incorporating quantum resources like memories or repeaters may also profoundly alter the underlying percolation landscape.
@@
These established mappings are, however, exclusively confined to discrete-variable (DV) systems (e.g., qubits). They all overlook an alternative architecture particularly prominent in optical settings: the continuous-variable (CV) system, leveraging the continuous degrees of freedom of light fields (amplitude and phase) [33, *Wang2007, *CSRNTJS2012]. In contrast to DV encoding, which is inherently reliant on random single-photon sources, the CV encoding is more advantageous with its unconditional, consistent generation of entanglement via nonlinear optical interactions, bypassing the hurdles of unpredictability in DV sources. This nonrandom nature unlocks considerable potential for the scalability of quantum optical applications [36, *JT2017, *FW2018, *JMNT2020], a potential now being realized with the advent of CV photonic chips [40].
@@
To address these conceptual gaps, here we introduce a Gaussian-to-Gaussian (G-G) DET scheme, wherein both the inputs and outputs are two-mode squeezed vacuum states (TMSVSs), defined as $|\psi^{r}\rangle=\sqrt{1-\tanh^{2}r}\sum\nolimits_{n=0}^{\infty}\tanh^{n}r|nn\rangle$, which are a class of entangled CV Gaussian states characterized by the squeezing parameter $r\in[0,\infty)$ [33, *Wang2007, *CSRNTJS2012]. Akin to DV-based DET [27], the G-G DET employs deterministic entanglement swapping [41] and concentration [42] adapted for TMSVSs, offering an avenue for entanglement distribution in optical systems.
@@
Further, we show that the NegPT exhibits a mixed-order phase transition [45], characterized by both a discontinuous change at a critical threshold $\chi_{\text{th}}$ and a divergent correlation length $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-z\nu}$. While mixed-order phase transitions have been widely observed in spin systems [45], interdependent networks [46, *AD2022], superconducting devices [48], overload systems [49], $k$-core percolation [50], and colloidal crystals [51], their observation in QNs is unprecedented. Our discovery extends the reach of mixed-order phase transitions beyond traditional statistical physics systems, into the rapidly evolving field of quantum information for the first time. Indeed, our finding of the NegPT, with a unique thermal critical exponent $z\nu\approx 1/2$ in Bethe lattices, clearly highlights a universality class that is distinct from both classical and concurrence percolation (where $z\nu\approx 1$[28, *OXSGBJ2022]. We also notice that the onset of mixed-order transitions may persist for generalized CV operations and even non-Gaussian states, suggesting that the NegPT may underlie the critical phenomena for generic CV-based QNs.
@@
Moreover, our theoretical analysis reveals a potential practical vulnerability within large-scale CV-based QNs: We show that the sharp transition at $\chi_{\text{th}}$ challenges stabilization via standard feedback, risking unstable “on/off” oscillations near the threshold—in stark contrast with continuous transitions. Given the prevalence of quantum feedback control [52, *quantum_control1999, *feedback_coherenct2000, *feedback2002, *feedback2002_2, *feedback2005, *quantum_control2008Domenico, *quantum_control2010, *feedback2012, *feedback2012_2, *feedback2017, *feedback_remote_entanglement2015] in quantum optical implementations, this instability acts as a crucial alert: scaling CV systems requires careful feedback strategies to maintain robust operation near the “brink of failure.” We initiate the characterization of this behavior and its network implications.
@@
Gaussian QN.—To begin with, we consider a Gaussian QN built by bipartite, pure Gaussian states $|\Psi\rangle_{AB}$ involving multiple modes (say, $K_{A}$ modes + $K_{B}$ modes) in general. A key insight for Gaussian states is that $|\Psi\rangle_{AB}$ can be mapped into a tensor product of $K\leq\min{(K_{A},K_{B})}$ TMSVSs (along with trivial vacuum states) via local Gaussian unitaries [64]: $|\Psi\rangle_{AB}=\bigotimes_{k=1}^{K}|\psi^{r_{k}}\rangle\otimes|00\dots\rangle$, where each $r_{k}$ corresponds to the squeezing parameter of the $k$-th TMSVS. This mode decomposition [64] simplifies our QN model to primarily focus on TMSVSs, such that each TMSVS represents a QN link between two nodes. A bipartite multi-mode pure Gaussian state can be treated as multi-links that connect the same pair of nodes multiple times, which are allowed in the formation of the QN.
@@
A DET scheme for series-parallel Gaussian QNs.—Our G-G DET scheme, following and extending the original DET scheme proposed for DV systems [27], consists of entanglement swapping and concentration operations applied to Gaussian states. The two operations, taking TMSVSs as the input and a new TMSVS as the output, are inherently deterministic in the CV domain. These operations can be intuitively understood as “series and parallel rules,” as commonly used in studying the effective resistance in circuit networks [27]. Hence, our G-G DET scheme can be applied to any network with a series-parallel topology [65]. The corresponding series (swapping) and parallel (concentration) rules are introduced as follows:
Table 2: G-G DET scheme. The scheme is composed of series and parallel rules derived from deterministic entanglement swapping and concentration protocols, respectively.
            LOCC Squeezing parameters $r_{1}>r_{2}>\dots$
Swapping (series) $\tanh{r}=\tanh{r_{1}}\tanh{r_{2}}\tanh{r_{3}}\dots$
Concentration (parallel) $\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}\cosh{r_{3}}\dots$
@@
  Exact rules         Approximate rules (by the star-mesh transform [28, *OXSGBJ2022])
Topology: Series Parallel Wheatstone bridge Kelvin bridge (higher-order topologies)
@@
$\mathrm{X}_{\text{SC}}$ $=\chi^{N}$ $={\chi}\left[\chi^{2}+\left(1-\chi^{2}\right)^{K}\right]^{-1/2}$ $\approx b(\chi,y(\chi))$ $\approx b(\chi,y(\chi\delta(\chi)))$dots
@@
$b(\chi,y)=\frac{\chi^{2}}{\sqrt{\chi^{4}+\left(1-y^{2}\right)\left[\left(\chi^{2}y^{2}-y^{2}+1\right)^{2}-\chi^{4}\right]}}$; $y(x)$ and $\delta(x)$ are the Y-$\Delta$ and $\Delta$-Y transforms of $x$, respectively (SM, Sec. II).
(1) G-G entanglement swapping.—Consider a setup involving three spatially separated parties: the Source ($S$), Relay ($R$), and Target ($T$). A TMSVS $|\psi^{r_{1}}\rangle$ is shared between $S$ and $R$, while another $|\psi^{r_{2}}\rangle$ is shared between $R$ and $T$. The two parties $S$ and $T$ do not directly share any entangled state. Nevertheless, executing homodyne measurement at $R$ and applying matching displacements at $S$ and $T$ can deterministically convert $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into a new TMSVS, $|\psi^{r}\rangle$, directly shared between $S$ and $T$ [41]. The new squeezing parameter $r$ is (ideally) given by $\tanh{r}=\tanh{r_{1}}\tanh{r_{2}}$ [41]. This can be extended to a one-dimensional chain consisting of a series of $N$ links corresponding to TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\cdots,|\psi^{r_{N}}\rangle$ (see Supplemental Materials (SM), Sec. A), respectively, leading to the series rule [Table. 2] for $|\psi^{r}\rangle$ being the final created TMSVS.
@@
(2) G-G entanglement concentration.—It is known that Gaussian states cannot be feasibly concentrated using standard optical components [66]. The only exception we know of is provided in Ref. [42], offering a conceptual construction of LOCC for concentrating entanglement from pure Gaussian states. Such LOCC must involve non-standard optical components. Assuming this, if we consider a tensor product of two TMSVSs shared by two parties $S$ and $T$, $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ ($r_{1}\geqslant r_{2}$), this state can be deterministically converted to a single TMSVS and a vacuum state, $|\psi^{r}\rangle\otimes|00\rangle$ by LOCC [42], where $\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}$ [42]. The result also scales to $K$ parallel links, $|\Psi\rangle_{{ST}}=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ (SM, Sec. B). By iteratively applying the Gaussian concentration $K-1$ times, the multi-mode state $|\Psi\rangle_{{ST}}$ is concentrated into a single TMSVS, leading to the parallel rule (Table 2).
@@
The parallel rule is not commutative due to the presence of a $\sinh r_{1}$ term before other $\cosh r_{k}$ terms. We find that the optimal entanglement of $|\psi^{r}\rangle$ is achieved by placing the largest squeezing parameter $r_{1}$ into $\sinh r_{1}$ (SM, Sec. B).
@@
Negativity percolation theory (NegPT).—The G-G DET scheme provides the foundation for a percolation-like theory of “quantum connectivity,” which we term negativity percolation. We consider a homogeneous QN where each link represents a TMSVS with an identical entanglement weight $\chi$. This framework is analogous to classical bond percolation on a network where each link is active with an identical probability $p$ [67].
@@
In classical percolation theory, the connectivity between two arbitrary sets of nodes, $S$ and $T$, can be quantified by the “sponge-crossing” probability, $P_{\text{SC}}$ [28, *OXSGBJ2022]. This quantity is the combination of probabilities over all paths connecting the two sets. Drawing a parallel, we introduce the “sponge-crossing” ratio negativity, $\mathrm{X}_{\text{SC}}$, which represents an analogous combination of all paths but is framed in terms of the ratio negativity $\chi$. For both classical and negativity percolation, this aggregation of all path contributions is calculated as follows:
@@
(1) When the network topology between $S$ and $T$ is series-parallel, $\mathrm{X}_{\text{SC}}$ can be decomposed into the iteration of the exact series and parallel rules. Notably, in a series network, $\mathrm{X}_{\text{SC}}=\chi^{N}$ has the same form as the probability $p$ or concurrence $c$ measures in classical or concurrence percolation [28, *OXSGBJ2022]. This similarity prompts us to refer to the theory as the NegPT, emphasizing on the special role that ratio negativity $\chi$ plays (analogous to $p$ or $c$).
@@
(2) When the network topology between $S$ and $T$ is not series-parallel, we do not know the exact transmission rules (beyond the scope of G-G DET). However, we can approximate these higher-order rules using transforms based only on the series/parallel rules (SM, Sec. II).
@@
Figure 1: Bethe lattice. 1 A Bethe lattice of degree $k$ (i.e., each node is incident to $k$ links) and network depth $l$ (the path length from the yellow node to the red nodes). 1 The sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ between $S$ and $T$ for various $k$ (right panel), satisfying the power law $\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\sim|\chi-\chi_{\text{th}}|^{1/2}$ as $\chi\to\chi_{\text{th}}^{+}$ (left panel). 1 When below the critical threshold ($\chi<\chi_{\text{th}}$), $\mathrm{X}_{\text{SC}}$ exhibits a plateau behavior until the network depth $l$ exceeds the correlation length $l^{*}$ (defined as the depth $l$ at which $\mathrm{X}_{\text{SC}}=0.5$), after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. 1 Near the critical threshold, we observe $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-0.508(9)}$, indicating $z\nu\approx 1/2$.
Mixed-order phase transition in CV-based QNs.—We focus on the Bethe lattice [67], an infinite tree-like series-parallel network where every node has an identical degree $k>2$ [Fig. 1]. The root (yellow) corresponds to the set $S$, and all the nodes (red) situated at the outermost boundary are collectively regarded as the set $T$. We define the (finite) network depth $l$ as the number of links in the shortest path between $S$ and $T$ (SM, Sec. A).
In the $l\to\infty$ limit, the Bethe lattice manifests self-similarity: the dark and light orange shaded regions are identical, as are the dark and light green shaded regions. This enables us to derive the exact self-consistent renormalization group [68, *AE2002] equation (SM, Sec. B). We observe a distinctive phenomenon: as $\chi$ increases from zero, $\mathrm{X}_{\text{SC}}$ initially remains zero until $\chi=\chi_{\text{th}}$, at which point $\mathrm{X}_{\text{SC}}$ exhibits an abrupt, discontinuous jump to a positive value $\mathrm{X}_{\text{SC}}^{+}$ before continuing to increase [Fig. 1]. For general $k$, we find
@@
where $T(k)=\left(k-1\right)^{-1/{\left(k-2\right)}}$ (SM, Sec. C). This is in stark contrast with classical or concurrence percolation, where the phase transition is shown to be continuous [28, *OXSGBJ2022]. Another difference is that in concurrence percolation, a nontrivial saturation point exists, above which the sponge-crossing entanglement becomes unity (even when the entanglement per link is below one) [28, *OXSGBJ2022]. Yet, there is no saturation point in the NegPT. Instead, we find $1-\mathrm{X}_{\text{SC}}\sim(1-\chi)^{k}$ when $\chi\to 1^{-}$, and $\mathrm{X}_{\text{SC}}$ becomes one if and only if $\chi=1$ (SM, Sec. D).

### v2

# Negativity Percolation in Continuous-Variable Quantum Networks

@@ Abstract
Quantum networks (QNs) have been predominantly driven by discrete‑variable (DV) architectures. Yet, optical platforms naturally generate Gaussian states—the common states of continuous-variable (CV) systems, making CV-based QNs an attractive route toward scalable, chip‑integrated quantum computation and communication. To bridge the gap between well‑studied DV entanglement percolation theories and their CV counterpart, we introduce a Gaussian-to-Gaussian entanglement distribution scheme that deterministically transports two‑mode squeezed vacuum states across large CV networks. Analysis of the scheme’s collective behavior using statistical-physics methods reveals a new form of entanglement percolation—negativity percolation theory (NegPT)—characterized by a bounded entanglement measure called the ratio negativity. We discover that NegPT exhibits a mixed-order phase transition, marked simultaneously by both an abrupt change in global entanglement and a long-range correlation between nodes. This distinctive behavior places CV-based QNs in a new universality class, fundamentally distinct from DV systems. Additionally, the abruptness of this transition introduces a critical vulnerability of CV-based QNs: conventional feedback mechanism becomes inherently unstable near the threshold, highlighting practical implications for stabilizing large-scale CV-based QNs. Our results unify statistical models for CV-based entanglement distribution and uncover previously unexplored critical phenomena unique to CV systems, providing valuable insights and guidelines essential for developing robust, feedback-stabilized QNs.
Introduction
Table 1: Comparison of percolation theories (PTs) for entanglement distribution on discrete-variable (DV) and continuous-variable (CV) quantum networks (QNs). The PTs are based on distinct entanglement measures: concurrence [49, 45] for DV qudit systems, and ratio negativity [97] for CV infinite-dimensional squeezed states.
  Entanglement distribution scheme PT / entanglement measure Phase transition type
@@ Introduction
DV Deterministic entanglement transmission (DET) ConPT / Concurrence $c$ Continuous (second order)
CV Gaussian-to-Gaussian (G-G) DET NegPT / Ratio negativity $\chi$ Discontinuous (mixed order)
@@
Quantum information technologies [74] fundamentally rely on our ability to distribute entanglement [10, 23], the process of establishing quantum correlations between distant nodes across network architectures, from chips to the internet. This capability empowers applications ranging from quantum computation [24, 50] to secure quantum communication [11, 18, 79, 72, 91, 93, 99]. The collective behavior of entanglement distribution across large-scale systems is often described by the quantum network (QN) model [1]. Within this framework, a QN is represented as a graph where nodes represent spatially separated parties, and links denote bipartite entangled states shared between adjacent nodes. The distribution process itself is executed via protocols based on local operations and classical communication (LOCC), such as entanglement swapping [100, 15] and entanglement concentration [9, 70], along with more sophisticated schemes [6, 19, 59, 75], applied across the network’s links to establish long-range correlations.
@@
Investigating how LOCC schemes distribute entanglement across a QN has uncovered fundamental correspondences to percolation theories. The ability to establish long-range entanglement in QN often mirrors the emergence of large-scale connectivity described by percolation theories—the statistical physics framework governing phase transitions on graphs [54, 21, 25, 13, 4, 12, 60]. This conceptual bridge between entanglement distribution and percolation was first built in 2007, linking a probabilistic entanglement distribution scheme [1] to classical bond percolation [35]. Later, the development of a more effective deterministic entanglement transmission (DET) scheme [65] (where LOCC steps succeed deterministically, eliminating classical randomness unlike earlier probabilistic schemes) led to its mapping onto an essentially distinct variant, concurrence percolation theory (ConPT) [66, 63]. This interplay continues with more recent mappings identified between quantum memories and continuum percolation [68], and between entanglement routing and path percolation [55, 67], underscoring how incorporating quantum resources like memories or repeaters may also profoundly alter the underlying percolation landscape.
@@
These established mappings are, however, exclusively confined to discrete-variable (DV) systems (e.g., qubits). They all overlook an alternative architecture particularly prominent in optical settings: the continuous-variable (CV) system, leveraging the continuous degrees of freedom of light fields (amplitude and phase) [17, 84, 85, 81]. In contrast to DV encoding, which is inherently reliant on random single-photon sources, the CV encoding is more advantageous with its unconditional, consistent generation of entanglement via nonlinear optical interactions, bypassing the hurdles of unpredictability in DV sources. This nonrandom nature unlocks considerable potential for the scalability of quantum optical applications [20, 29, 41, 30], a potential now being realized with the advent of CV photonic chips [53].
@@
To address these conceptual gaps, here we introduce a Gaussian-to-Gaussian (G-G) DET scheme, wherein both the inputs and outputs are two-mode squeezed vacuum states (TMSVSs), defined as $|\psi^{r}\rangle=\sqrt{1-\tanh^{2}r}\sum\nolimits_{n=0}^{\infty}\tanh^{n}r|nn\rangle$, which are a class of entangled CV Gaussian states characterized by the squeezing parameter $r\in[0,\infty)$ [17, 84, 85]. Akin to DV-based DET [65], the G-G DET employs deterministic entanglement swapping [82] and concentration [51] adapted for TMSVSs, offering an avenue for entanglement distribution in optical systems.
@@
Further, we show that the NegPT exhibits a mixed-order phase transition [71], characterized by both a discontinuous change at a critical threshold $\chi_{\text{th}}$ and a divergent correlation length $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-z\nu}$. While mixed-order phase transitions have been widely observed in spin systems [71], interdependent networks [98], superconducting devices [14], overload systems [77], $k$-core percolation [92], and colloidal crystals [3], their observation in QNs is unprecedented. Our discovery extends the reach of mixed-order phase transitions beyond traditional statistical physics systems, into the rapidly evolving field of quantum information for the first time. Indeed, our finding of the NegPT, with a unique thermal critical exponent $z\nu\approx 1/2$ in Bethe lattices, clearly highlights a universality class that is distinct from both classical and concurrence percolation (where $z\nu\approx 1$) [66, 63]. We also notice that the onset of mixed-order transitions may persist for generalized CV operations and even non-Gaussian states, suggesting that the NegPT may underlie the critical phenomena for generic CV-based QNs.
@@
Moreover, our theoretical analysis reveals a potential practical vulnerability within large-scale CV-based QNs: We show that the sharp transition at $\chi_{\text{th}}$ challenges stabilization via standard feedback, risking unstable “on/off” oscillations near the threshold—in stark contrast with continuous transitions. Given the prevalence of quantum feedback control [88, 31, 61, 62, 47, 52, 26, 32, 95, 80, 96, 64] in quantum optical implementations, this instability acts as a crucial alert: scaling CV systems requires careful feedback strategies to maintain robust operation near the “brink of failure.” We initiate the characterization of this behavior and its network implications.
Results
.1 Gaussian QN
@@ Results
To begin with, we consider a Gaussian QN built by bipartite, pure Gaussian states $|\Psi\rangle_{AB}$ involving multiple modes (say, $K_{A}$ modes + $K_{B}$ modes) in general. A key insight for Gaussian states is that $|\Psi\rangle_{AB}$ can be mapped into a tensor product of $K{\color[rgb]{0,0,0}\leq}\min{(K_{A},K_{B})}$ TMSVSs (along with trivial vacuum states) via local Gaussian unitaries [16]: $|\Psi\rangle_{AB}=\bigotimes_{k=1}^{K}|\psi^{r_{k}}\rangle\otimes|00\dots\rangle$, where each $r_{k}$ corresponds to the squeezing parameter of the $k$-th TMSVS. This mode decomposition [16] simplifies our QN model to primarily focus on TMSVSs, such that each TMSVS represents a QN link between two nodes. A bipartite multi-mode pure Gaussian state can be treated as multi-links that connect the same pair of nodes multiple times, which are allowed in the formation of the QN.
Figure 1: Gaussian-to-Gaussian deterministic entanglement transmission (G-G DET) scheme. Applicable to Gaussian quantum networks (QN), the scheme consists of two LOCC protocols: (a) Entanglement swapping, facilitated by homodyne detection and displacement [82]; (b) Entanglement concentration, facilitated by non-standard optical components [51]. Both protocols are deterministic, taking two (or more) TMSVS $|\psi^{r_{i}}\rangle$ as input and a new TMSVS $|\psi^{r}\rangle$ as output. (c) The two LOCC protocols map to series and parallel rules, respectively, to construct G-G DET. (d) Consider a QN example built upon three node. The G-G DET scheme consists of two steps: First, the parallel rule converts the states $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ ($r_{1}\geq r_{2}$) into $|\psi^{r_{1,2}}\rangle$ with $\sinh r_{1,2}=\sinh r_{1}\cosh r_{2}$ between $S$ and $R$; second, the series rule transforms $|\psi^{r_{1,2}}\rangle$ and $|\psi^{r_{3}}\rangle$ to the final state $|\psi^{r}\rangle$ with the ratio negativity $\text{X}_{\text{SC}}=\tanh r_{1,2}\tanh r_{3}$ between $S$ and $T$.
.2 A DET scheme for series-parallel Gaussian QNs
@@
Our G-G DET scheme, following and extending the original DET scheme proposed for DV systems [65], consists of entanglement swapping and concentration operations applied to Gaussian states. The two operations, taking TMSVSs as the input and a new TMSVS as the output, are inherently deterministic in the CV domain. These operations can be intuitively understood as “series and parallel rules,” as commonly used in studying the effective resistance in circuit networks [65]. Hence, our G-G DET scheme can be applied to any network with a series-parallel topology [33]. The corresponding series (swapping) and parallel (concentration) rules are introduced as follows:
First, we introduce the G-G entanglement swapping. Consider a setup involving three spatially separated parties: the Source ($S$), Relay ($R$), and Target ($T$). A TMSVS $|\psi^{r_{1}}\rangle$ is shared between $S$ and $R$, while another $|\psi^{r_{2}}\rangle$ is shared between $R$ and $T$ [Fig. 1(a)]. The two parties $S$ and $T$ do not directly share any entangled state. Nevertheless, executing homodyne measurement at $R$ and applying matching displacements at $S$ and $T$ can deterministically convert $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into a new TMSVS, $|\psi^{r}\rangle$, directly shared between $S$ and $T$ [82]. The new squeezing parameter $r$ is (ideally) given by $\tanh{r}=\tanh{r_{1}}\tanh{r_{2}}$ [82]. This can be extended to a one-dimensional chain consisting of a series of $N$ links corresponding to TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\cdots,|\psi^{r_{N}}\rangle$ (see Methods), respectively, leading to the series rule [Fig. 1(c)] for $|\psi^{r}\rangle$ being the final created TMSVS.
@@
  Exact rules         Approximate rules (by the star-mesh transform [66, 63])
Topology: Series Parallel Wheatstone bridgea Kelvin bridgea (higher-order topologies)
@@
$\mathrm{X}_{\text{SC}}$ $\chi^{N} =$${\chi}\left[\chi^{2}+\left(1-\chi^{2}\right)^{K}\right]^{-1/2}$ $\approx b(\chi,y(\chi))$ $\approx b(\chi,y(\chi\delta(\chi)))$ $\dots$
  • a
@@
    $b(\chi,y)=\frac{\chi^{2}}{\sqrt{\chi^{4}+\left(1-y^{2}\right)\left[\left(\chi^{2}y^{2}-y^{2}+1\right)^{2}-\chi^{4}\right]}}$; $y(x)$ and $\delta(x)$ are the Y-$\Delta$ and $\Delta$-Y transforms of $x$, respectively (see Methods).
@@
We then give the G-G entanglement concentration. It is known that Gaussian states cannot be feasibly concentrated using standard optical components [34]. The only exception we know of is provided in Ref. [51], offering a conceptual construction of LOCC for concentrating entanglement from pure Gaussian states. Such LOCC must involve non-standard optical components. Assuming this, if we consider a tensor product of two TMSVSs shared by two parties $S$ and $T$, $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ ($r_{1}\geq r_{2}$), this state can be deterministically converted to a single TMSVS and a vacuum state, $|\psi^{r}\rangle\otimes|00\rangle$ by LOCC [Fig. 1(b)] [51], where $\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}$ [51]. The result also scales to $K$ parallel links, $|\Psi\rangle_{{ST}}=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ (see Methods). By iteratively applying the Gaussian concentration K1K-1 times, the multi-mode state $|\Psi\rangle_{{ST}}$ is concentrated into a single TMSVS, leading to the parallel rule [Fig. 1(c)].
@@
The parallel rule is not commutative due to the presence of a $\sinh r_{1}$ term before other $\cosh r_{k}$ terms. We find that the optimal entanglement of $|\psi^{r}\rangle$ is achieved by placing the largest squeezing parameter $r_{1}$ into $\sinh r_{1}$ (see Methods).
.3 Negativity percolation theory (NegPT)
@@
The G-G DET scheme provides the foundation for a percolation-like theory of “quantum connectivity,” which we term negativity percolation. We consider a homogeneous QN where each link represents a TMSVS with an identical entanglement weight $\chi$. This framework is analogous to classical bond percolation on a network where each link is active with an identical probability $p$ [8].
@@
In classical percolation theory, the connectivity between two arbitrary sets of nodes, $S$ and $T$, can be quantified by the “sponge-crossing” probability, $P_{\text{SC}}$ [66, 63]. This quantity is the combination of probabilities over all paths connecting the two sets. Drawing a parallel, we introduce the “sponge-crossing” ratio negativity, $\mathrm{X}_{\text{SC}}$, which represents an analogous combination of all paths but is framed in terms of the ratio negativity. Specifically, consider a Gaussian QN with endpoints $S$ and $T$ that initially share no entanglement. By implementing the G-G DET scheme across the QN, $S$ and $T$ ultimately share a TMSVS $|\psi^{r}\rangle$, with entanglement established between them. The final entanglement is quantified by the ratio negativity of $|\psi^{r}\rangle$, which we refer to as the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$.
For both classical and negativity percolation, this aggregation of all path contributions is calculated as follows:
@@
(1) When the network topology between $S$ and $T$ is series-parallel, $\mathrm{X}_{\text{SC}}$ can be decomposed into the iteration of the exact series and parallel rules [e.g., Fig. 1(d)]. Additionally, in a series network of $N$ TMSVSs with identical ratio negativity $\chi$, $\mathrm{X}_{\text{SC}}=\chi^{N}$ has the same form as the probability measure $p$ (or concurrence $c$) in classical (or concurrence) percolation [66, 63]. This similarity prompts us to refer to the theory as the NegPT, emphasizing on the special role that ratio negativity $\chi$ plays that is analogous to $p$ (or $c$).
@@
(2) When the network topology between $S$ and $T$ is not series-parallel, we do not know the exact transmission rules (beyond the scope of G-G DET). However, we can approximate these higher-order rules using transforms based only on the series/parallel rules (see Methods).
@@
Figure 2: Bethe lattice. (a) A Bethe lattice of degree $k$ (i.e., each node is incident to $k$ links) and network depth $l$ (the path length from the yellow node to the red nodes). (b) The sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ between $S$ and $T$ for various $k$ (right panel), satisfying the power law $\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\sim|\chi-\chi_{\text{th}}|^{0.47(5)}$ as $\chi\to\chi_{\text{th}}^{+}$ (left panel). The numerical value $0.47\pm 0.05$ is derived by a linear least-squares fit to the sixteen data points. (c) When $\chi\to\chi_{\text{th}}^{-}$, $\mathrm{X}_{\text{SC}}$ exhibits a plateau behavior until the network depth $l$ exceeds the correlation length $l^{*}$ (defined as the depth $l$ at which $\mathrm{X}_{\text{SC}}=0.5$), after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. (d) Near the critical threshold, we observe $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-0.508(9)}$, indicating $z\nu\approx 1/2$.
.4 Mixed-order phase transition
We start by focusing on the Bethe lattice [8], an infinite tree-like series-parallel network where every node has an identical degree $k>2$ [Fig. 2(a)]. The root (yellow) corresponds to the set $S$, and all the nodes (red) situated at the outermost boundary are collectively regarded as the set $T$. We define the (finite) network depth $l$ as the number of links in the shortest path between $S$ and $T$ (see Supplemental Information (SI), Sec. I A).
In the $l\to\infty$ limit, the Bethe lattice manifests self-similarity: the dark and light orange shaded regions are identical, as are the dark and light green shaded regions. This enables us to derive the exact self-consistent renormalization group [22, 76] equation (SI, Sec. I B). We observe a distinctive phenomenon: as $\chi$ increases from zero, $\mathrm{X}_{\text{SC}}$ initially remains zero until $\chi=\chi_{\text{th}}$, at which point $\mathrm{X}_{\text{SC}}$ exhibits an abrupt, discontinuous jump to a positive value $\mathrm{X}_{\text{SC}}^{+}$ before continuing to increase [Fig. 2(b)]. For general $k$, we find
@@
where $T(k)=\left(k-1\right)^{-1/{\left(k-2\right)}}$ (SI, Sec. I C). This is in stark contrast with classical or concurrence percolation, where the phase transition is shown to be continuous [66, 63]. Another difference is that in concurrence percolation, a nontrivial saturation point exists, above which the sponge-crossing entanglement becomes unity (even when the entanglement per link is below one) [66, 63]. Yet, there is no saturation point in the NegPT. Instead, we find $1-\mathrm{X}_{\text{SC}}\sim(1-\chi)^{k}$ when $\chi\to 1^{-}$, and $\mathrm{X}_{\text{SC}}$ becomes one if and only if $\chi=1$ (SI, Sec. I D).